Please help me solve this equation
-2x+3y-5z+w=0
4x-6y+10z-2w=0

Answer by checkley77(12844)

-2x+3y-5z+w=0
2X-3Y+5Z-W=0 SUBTRACT FROM THE OTHER EQUATION.
------------------
4x-6y+10z-2w=0
2(2X-3Y)+2(5Z-W)=0
-(2X-3Y)-(5Z-W)=0
---------------------------
(2X-3Y)+(5Z-W) ANS.